Determine whether each pair of lines is a parallel, perpendicular, or neither.
Line 1 goes through (-4,10) and (2,-3)
Line 2 goes through (-3,8) and (2,-5)

Answer by RAY100(1637)

(-4,10)(2,-3),,,,,m=(y2-y1)/(x2-x1),,,,,m=(10-(-3))/(-4-2),,,m=- 13/6
.
(-3,8)(2,-5),,,,m= (8-(-5)) / (-3-2),,,,m=-13 / 5
.
not parallel (m1 not = m2)
.
not perpendicular ( m1 not = -1/m2)
.
Neither, ok,,, intercept at (-2.615,7)